# Entropy factorization via curvature

### v1

@@ 1.1 Motivation
for all bounded measurable functions $g_{A}:\mathbb{R}^{|A^{c}|}\mapsto\mathbb{R}_{+}$, where, for all $A\subset[n]$, $c_{A}:=\frac{\theta_{A}}{1-\kappa}$, and $Z_{A}$ denotes the set of variables $(Z_{i},i\in A)$. We note that if $f$ is a probability density with respect to $\mathbf{P}$, so that $f\mathbf{P}$ defines a probability measure on $\Omega$, then ${\mathbf{E}}_{A}[f]$ is the density of the pushforward of $f\mathbf{P}$ under the projection $Z\mapsto Z_{A^{c}}$ (the marginal of $f\mathbf{P}$ on $Z_{A^{c}}$). In fact, the equivalence between (7) and (4) applies more generally when the projection $Z\mapsto Z_{A^{c}}$ is replaced by an arbitrary measurable map $Z\mapsto B_{A}(Z)$, and ${\mathbf{E}}_{A}[f]$ is replaced by the conditional expectation ${\mathbf{E}}[f|B_{A}(Z)]$; see [17, Theorem 2.1]. The inequalities (7) obtained in this way are referred to as Brascamp-Lieb (B-L) type inequality, in analogy with the classical B-L inequality, which corresponds to the setting where $\mathbf{P}$ is the Lebesgue measure on $\mathbb{R}^{n}$ and the measurable maps $B_{A}$ are linear; see [10, 26, 4, 7, 25].
@@
We adopt a more general viewpoint on the inequalities (7) and (4), by replacing conditional expectations ${\mathbf{E}}_{A}[f]$ with arbitrary, not necessarily self-adjoint, Markov operators $f\mapsto T_{A}f$. It turns out that it is not hard to extend the duality to the new framework, thus generalizing the equivalence proved in [17]; see Section 1.2. We hope that this point of view will provide a fruitful generalization of the classical duality, possibly allowing for interesting developments and new B-L type inequalities.
@@ 1.2 General framework
We remark that when the Markov operators $T_{i}$ are conditional expectations of the form $T_{i}f={\mathbf{E}}[f|\mathcal{F}_{i}]$ for some sub $\sigma$-algebras $\mathcal{F}_{i}$, Theorem 1 yields the duality proved in [17, Theorem 2.1]. In that case $T_{i}$ is an orthogonal projection in $L^{2}$, and therefore $T_{i}$ is self-adjoint, and by the chain rule (5) one has that both statements (1) and (2) are equivalent to an entropy factorization statement of the form
@@
We note that the case where $T_{i}$ is not a conditional expectation goes beyond the common B-L inequality setup and we are not aware of previous work addressing the duality in this generality. The proof, however, boils down to essentially the same argument as in [17], which is based on the well known variational principle for relative entropy stating that
@@ 1.4 Some examples
The discrete Gaussian free field is a Gaussian probability measure on $\mathbb{R}^{n}$, with covariance of the form $\Gamma=({\rm Id}-P)^{-1}$, where $P$ is a symmetric $n\times n$ sub-stochastic matrix, that is $P_{i,j}\geq 0$ for all $i,j\in[n]$, $\sum_{j}P_{i,j}\leq 1$ for all $i\in[n]$, and such that ${\rm Id}-P$ is invertible. We also assume that $P$ is irreducible, that is for all $i,j$ there exists $k\in\mathbb{N}$ such that $P^{k}_{i,j}>0$. A well studied example in equilibrium statistical mechanics is obtained by taking $P$ as the transition matrix of the random walk on $\mathbb{Z}^{d}$ that is killed upon exiting a finite domain $V\subset\mathbb{Z}^{d}$, see e.g. [23, Chapter 8].
@@
Combining Theorem 3 with Theorem 1 one obtains equivalent B-L inequalities for the gaussian measures. The latter were in fact already known, since they can be directly related to the classical B-L inequalities for Lebesgue measures, see e.g. [5, Fact 6]. A direct proof of the gaussian B-L inequality is also possible along the lines of [25], see [20, Theorem 3]. Thus, one could have obtained Theorem 3 from these known estimates. However, the purpose of Theorem 3 is to illustrate that an entirely different approach based on curvature can be used to obtain optimal entropy factorization estimates, and, in principle, it can yield an alternative proof of the classical B-L inequalities as well. One caveat is that we assume the specific form $\Gamma=({\rm Id}-P)^{-1}$ for the covariance matrix of our gaussian measure, whereas it would be desirable to have the result for all positive definite matrices $\Gamma$. For technical reasons, in this more general setup at the moment we can only prove the desired estimate up to a factor $2$, that is we can show that $\sigma\leq 2\kappa\leq 2\sigma$, for an explicit spectral quantity $\sigma=\sigma(\Gamma,\theta)$, see Remark 5 for more details. We believe it should be possible to obtain $\kappa=\lambda=\sigma$ for all $(\Gamma,\theta)$, as we prove in the case of $\Gamma=({\rm Id}-P)^{-1}$ with Glauber dynamics in Theorem 3, thus obtaining full applicability of our argument to derive Gaussian and classical B-L inequalities.
@@
and that the only extremal functions are Dirac masses. Here one has $\theta_{\star\star}=\frac{\ell(\ell-1)}{n(n-1)}$ which is smaller than $\kappa_{\ell,n}$ for $\ell\in\{2,\dots,n-1\}$, and therefore Theorem 4 is not optimal for such homogeneous weights. However, Theorem 4, as well as [5, Proposition 21], addresses the entropy factorization problem for arbitrary weights $(\theta_{A})$, and in this generality they may provide the best known bounds.
@@
Letting $\rho_{\rm MLS}$ denote the modified log-Sobolev constant from (28), the simple bound $\rho_{\rm MLS}\geq\kappa$ allows one to obtain $\rho_{\rm MLS}\geq\frac{2}{n(n-1)}$, which recovers (and actually improves by a factor 2) an estimate previously shown by Villani in [37, Theorem 6.1] by a different type of semigroup interpolation.
@@
For $p=2$ this coincides with the result in [5, Proposition 11]. A related estimate for all $p>0$, under additional symmetry assumptions on $f$, was obtained in [5, Corollary 16]. Note that the above statement is formally equivalent to Theorem 4. In fact, to prove the conjecture one may proceed exactly as in our proof of Theorem 4, namely by exhibiting a metric on $\Omega$ for which Assumptions (i) and (ii) in Theorem 2 are satisfied with $\kappa=\theta_{\star\star}$, when the Markov operators are given by the conditional expectations $T_{A}f={\mathbf{E}}_{A}f$. As we show in Section 3.3, this can be achieved by taking the special metric
@@ Acknowledgement.
We thank Thomas Courtade for helpful conversations on the B-L inequalities for Gaussian measures, as well as Francesco Pedrotti and Sam Power for many relevant comments and references. P.C. warmly thanks the research center CEREMADE at Paris Dauphine for the kind hospitality.
@@ 3.1 Gaussian Free Fields: Proof of Theorem 3
expresses $Z$ as the sum of a $\mathcal{G}-$measurable vector and a $\mathcal{G}-$independent one. The claim now follows from a standard property of conditional expectation (see, e.g. [39, Section 9.10]). ∎
@@
The second property in Remark 1 is the regularity $\mathrm{Lip}(P_{\tau}f)<\infty$ for all $f\in L^{\infty}$, $\tau>0$. Since our metric is comparable to Euclidean norm, this follows from the well known regularizing property of the Ornstein-Uhlenbeck semi-group, see e.g. [2, Th. 4.7.2].
@@
Since $21-\sqrt{1-\sigma}\geq\sigma/$ 2 this estimate is sharp up to a factor at most $2$. As mentioned in Remark 4, it should be possible to improve this argument to obtain the sharp result $\kappa=\sigma$ for all positive definite $\Gamma$ and all probability vector $\theta$. The identity $\kappa=\sigma$ could be in fact obtained via Theorem 1 using known B-L inequalities for Gaussian measures, see [20, Theorem 3].
@@ 3.2 Permutations: proof of Theorem 4
To check the positive curvature, we use a finer coupling. We first observe that the metric $\,{\rm dist}(\sigma,\eta)$ is generated by pairs of configurations that differ by exactly one swap, and thus using a telescopic decomposition, using the so-called Gluing Lemma (see, e.g., [38, Lemma 7.6]), we may restrict to a pair $(\sigma,\eta)$ such that $\sigma_{\ell}=\eta_{\ell}$, for all $\ell\neq i,j$, and $(\sigma_{i},\sigma_{j})=(\eta_{j},\eta_{i})$, for a fixed pair of distinct indexes $i,j\in[n]$. This reduction step is also known as the path coupling lemma [12]. Given such a pair $(\sigma,\eta)$, the coupling of $(\sigma^{\prime},\eta^{\prime})$ is described as follows. Let $\pi$ be a uniformly random permutation of the labels $\{j\in[n]:j\in A\}$. If either $i$ or $j$ (or both) are in $A^{c}$ then we update as before by setting $(\sigma^{\prime},\eta^{\prime})=(\sigma\pi,\eta\pi)$. If instead $i,j$ are both in $A$, then we ensure coalescence by replacing $(\sigma,\eta)$ with $(\sigma\pi,\sigma\pi)$. This shows that
@@ 3.5 Langevin diffusion: proof of Theorem 6
We may now apply the case $M=1$ of Theorem 2 to the operator $T=P_{t}$, for a fixed $t>0$. Since $P_{t}^{\star}=P_{t}$ and $W_{1}\leq W_{\infty}$, Assumptions (i)-(ii) hold with $\ell=e^{-\rho t}$ and $\kappa=1-e^{-2\rho t}$. Moreover, the regularity property (iii) classically holds in this setting (see, e.g., [2, Th. 4.7.2]). The desired conclusion readily follows.

### v2

@@ 1.1 Motivation
for all bounded measurable functions $g_{A}:\mathbb{R}^{|A^{c}|}\mapsto\mathbb{R}_{+}$, where, for all $A\subset[n]$, $c_{A}:=\frac{\theta_{A}}{1-\kappa}$, and $Z_{A}$ denotes the set of variables $(Z_{i},i\in A)$. We note that if $f$ is a probability density with respect to $\mathbf{P}$, so that $f\mathbf{P}$ defines a probability measure on $\Omega$, then ${\mathbf{E}}_{A}[f]$ is the density of the pushforward of $f\mathbf{P}$ under the projection $Z\mapsto Z_{A^{c}}$ (the marginal of $f\mathbf{P}$ on $Z_{A^{c}}$). In fact, the equivalence between (7) and (4) applies more generally when the projection $Z\mapsto Z_{A^{c}}$ is replaced by an arbitrary measurable map $Z\mapsto B_{A}(Z)$, and ${\mathbf{E}}_{A}[f]$ is replaced by the conditional expectation ${\mathbf{E}}[f|B_{A}(Z)]$; see [18, Theorem 2.1]. The inequalities (7) obtained in this way are referred to as Brascamp-Lieb (B-L) type inequality, in analogy with the classical B-L inequality, which corresponds to the setting where $\mathbf{P}$ is the Lebesgue measure on $\mathbb{R}^{n}$ and the measurable maps $B_{A}$ are linear; see [11, 27, 5, 8, 26].
@@
We adopt a more general viewpoint on the inequalities (7) and (4), by replacing conditional expectations ${\mathbf{E}}_{A}[f]$ with arbitrary, not necessarily self-adjoint, Markov operators $f\mapsto T_{A}f$. It turns out that it is not hard to extend the duality to the new framework, thus generalizing the equivalence proved in [18]; see Section 1.2.
@@ 1.2 General framework
We remark that when the Markov operators $T_{i}$ are conditional expectations of the form $T_{i}f={\mathbf{E}}[f|\mathcal{F}_{i}]$ for some sub $\sigma$-algebras $\mathcal{F}_{i}$, Theorem 1 yields the duality proved in [18, Theorem 2.1]. In that case $T_{i}$ is an orthogonal projection in $L^{2}$, and therefore $T_{i}$ is self-adjoint, and by the chain rule (5) one has that both statements (1) and (2) are equivalent to an entropy factorization statement of the form
@@
We note that the case where $T_{i}$ is not a conditional expectation goes beyond the common B-L inequality setup. The proof, however, boils down to essentially the same argument as in [18], which is based on the well known variational principle for relative entropy stating that
@@
After completing this work, we learned that the generalized form of duality proposed in Theorem 1 had already been observed; see [1] for an early result in the case $M=1$, and [28] for a general result analogous to the one considered here. We also refer to [28] for a list of interesting applications of this generalized framework, such as hypercontractivity, strong data processing inequalities and other information-theoretic questions.
@@ 1.4 Some examples
The discrete Gaussian free field is a Gaussian probability measure on $\mathbb{R}^{n}$, with covariance of the form $\Gamma=({\rm Id}-P)^{-1}$, where $P$ is a symmetric $n\times n$ sub-stochastic matrix, that is $P_{i,j}\geq 0$ for all $i,j\in[n]$, $\sum_{j}P_{i,j}\leq 1$ for all $i\in[n]$, and such that ${\rm Id}-P$ is invertible. We also assume that $P$ is irreducible, that is for all $i,j$ there exists $k\in\mathbb{N}$ such that $P^{k}_{i,j}>0$. A well studied example in equilibrium statistical mechanics is obtained by taking $P$ as the transition matrix of the random walk on $\mathbb{Z}^{d}$ that is killed upon exiting a finite domain $V\subset\mathbb{Z}^{d}$, see e.g. [24, Chapter 8].
@@
Combining Theorem 3 with Theorem 1 one obtains equivalent B-L inequalities for the gaussian measures. The latter were in fact already known, since they can be directly related to the classical B-L inequalities for Lebesgue measures, see e.g. [6, Fact 6]. A direct proof of the gaussian B-L inequality is also possible along the lines of [26], see [21, Theorem 3]. Thus, one could have obtained Theorem 3 from these known estimates. However, the purpose of Theorem 3 is to illustrate that an entirely different approach based on curvature can be used to obtain optimal entropy factorization estimates, and, in principle, it can yield an alternative proof of the classical B-L inequalities as well. One caveat is that we assume the specific form $\Gamma=({\rm Id}-P)^{-1}$ for the covariance matrix of our gaussian measure, whereas it would be desirable to have the result for all positive definite matrices $\Gamma$. For technical reasons, in this more general setup at the moment we can only prove the desired estimate up to a factor $2$, that is we can show that $\sigma\leq 2\kappa\leq 2\sigma$, for an explicit spectral quantity $\sigma=\sigma(\Gamma,\theta)$, see Remark 5 for more details. We believe it should be possible to obtain $\kappa=\lambda=\sigma$ for all $(\Gamma,\theta)$, as we prove in the case of $\Gamma=({\rm Id}-P)^{-1}$ with Glauber dynamics in Theorem 3, thus obtaining full applicability of our argument to derive Gaussian and classical B-L inequalities.
@@
and that the only extremal functions are Dirac masses. Here one has $\theta_{\star\star}=\frac{\ell(\ell-1)}{n(n-1)}$ which is smaller than $\kappa_{\ell,n}$ for $\ell\in\{2,\dots,n-1\}$, and therefore Theorem 4 is not optimal for such homogeneous weights. However, Theorem 4, as well as [6, Proposition 21], addresses the entropy factorization problem for arbitrary weights $(\theta_{A})$, and in this generality they may provide the best known bounds.
@@
Letting $\rho_{\rm MLS}$ denote the modified log-Sobolev constant from (28), the simple bound $\rho_{\rm MLS}\geq\kappa$ allows one to obtain $\rho_{\rm MLS}\geq\frac{2}{n(n-1)}$, which recovers (and actually improves by a factor 2) an estimate previously shown by Villani in [39, Theorem 6.1] by a different type of semigroup interpolation.
@@
For $p=2$ this coincides with the result in [6, Proposition 11]. A related estimate for all $p>0$, under additional symmetry assumptions on $f$, was obtained in [6, Corollary 16]. Note that the above statement is formally equivalent to Theorem 4. In fact, to prove the conjecture one may proceed exactly as in our proof of Theorem 4, namely by exhibiting a metric on $\Omega$ for which Assumptions (i) and (ii) in Theorem 2 are satisfied with $\kappa=\theta_{\star\star}$, when the Markov operators are given by the conditional expectations $T_{A}f={\mathbf{E}}_{A}f$. As we show in Section 3.3, this can be achieved by taking the special metric
@@ Acknowledgement.
We thank Thomas Courtade for helpful conversations on the B-L inequalities, as well as Francesco Pedrotti and Sam Power for many relevant comments and references. P.C. warmly thanks the research center CEREMADE at Paris Dauphine for the kind hospitality.
@@ 3.1 Gaussian Free Fields: Proof of Theorem 3
expresses $Z$ as the sum of a $\mathcal{G}-$measurable vector and a $\mathcal{G}-$independent one. The claim now follows from a standard property of conditional expectation (see, e.g. [41, Section 9.10]). ∎
@@
The second property in Remark 1 is the regularity $\mathrm{Lip}(P_{\tau}f)<\infty$ for all $f\in L^{\infty}$, $\tau>0$. Since our metric is comparable to Euclidean norm, this follows from the well known regularizing property of the Ornstein-Uhlenbeck semi-group, see e.g. [3, Th. 4.7.2].
@@
Since $21-\sqrt{1-\sigma}\geq\sigma/$ 2 this estimate is sharp up to a factor at most $2$. As mentioned in Remark 4, it should be possible to improve this argument to obtain the sharp result $\kappa=\sigma$ for all positive definite $\Gamma$ and all probability vector $\theta$. The identity $\kappa=\sigma$ could be in fact obtained via Theorem 1 using known B-L inequalities for Gaussian measures, see [21, Theorem 3].
@@ 3.2 Permutations: proof of Theorem 4
To check the positive curvature, we use a finer coupling. We first observe that the metric $\,{\rm dist}(\sigma,\eta)$ is generated by pairs of configurations that differ by exactly one swap, and thus using a telescopic decomposition, using the so-called Gluing Lemma (see, e.g., [40, Lemma 7.6]), we may restrict to a pair $(\sigma,\eta)$ such that $\sigma_{\ell}=\eta_{\ell}$, for all $\ell\neq i,j$, and $(\sigma_{i},\sigma_{j})=(\eta_{j},\eta_{i})$, for a fixed pair of distinct indexes $i,j\in[n]$. This reduction step is also known as the path coupling lemma [13]. Given such a pair $(\sigma,\eta)$, the coupling of $(\sigma^{\prime},\eta^{\prime})$ is described as follows. Let $\pi$ be a uniformly random permutation of the labels $\{j\in[n]:j\in A\}$. If either $i$ or $j$ (or both) are in $A^{c}$ then we update as before by setting $(\sigma^{\prime},\eta^{\prime})=(\sigma\pi,\eta\pi)$. If instead $i,j$ are both in $A$, then we ensure coalescence by replacing $(\sigma,\eta)$ with $(\sigma\pi,\sigma\pi)$. This shows that
@@ 3.5 Langevin diffusion: proof of Theorem 6
We may now apply the case $M=1$ of Theorem 2 to the operator $T=P_{t}$, for a fixed $t>0$. Since $P_{t}^{\star}=P_{t}$ and $W_{1}\leq W_{\infty}$, Assumptions (i)-(ii) hold with $\ell=e^{-\rho t}$ and $\kappa=1-e^{-2\rho t}$. Moreover, the regularity property (iii) classically holds in this setting (see, e.g., [3, Th. 4.7.2]). The desired conclusion readily follows.